OpenStudy (anonymous):

2(x+3)+3(x-5)=8(x-6) <--Need Help

OpenStudy (anonymous):

2x + 6 + 3x - 15 = 8x - 48 6 - 15 + 48 = 8x -2x - 3x 39 = 3x x = 13